v= 3i + (4-10t)j
Find the equation of trajectory.
1
Expert's answer
2021-08-25T16:58:38-0400

"V=3i+(4-10t)j"

"dx\/dt=3"

"dx=3dt"

"x=3t" , "t=x\/3"


now

"dy\/dt=4-10t"

"dy=(4-10t)dt"

"y=4t-\frac{10t^2}{2}=4t-5t^2"

"t=x\/3"

"y=4x\/3-5(x\/3)^2"

"y=\frac{4x}{3}-\frac{5x^2}{9}"